# GUP-modified $\alpha$-attractor inflation in light of ACT observations

### v1

@@ Abstract
Here, the $\alpha$-attractor inflation is investigated within a framework incorporating a minimal measurable length, as implemented by the Generalized Uncertainty Principle (GUP). The GUP modifications to the Friedmann equations and cosmological perturbation parameters are employed to assess the model observational viability against recent high-precision data, primarily from the Atacama Cosmology Telescope Data Release 6 (ACT DR6) combined with Planck, DESI BAO, and BICEP/Keck 2018 observations (P-ACT-LB-BK18). Our results indicate that while the standard $\alpha$-attractor model ($\beta=0$) for $N_{*}=60$ lies near the $2\sigma$ boundary of the P-ACT-LB-BK18 dataset, the inclusion of GUP effects significantly shifts the predictions for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. Notably, for the GUP parameter $\beta$ values surpassing approximately $10^{13}$, the model predictions are shown to fall well within the 68% CL region of this combined observational data, enhancing its consistency. The viable parameter space in the $\alpha-\beta$ plane is explored, revealing that a maximum $\alpha\simeq 0.74$ is consistent with the 68% CL for $\beta\simeq 1.4\times 10^{13}$. In addition, for smaller values of $\alpha$, the predictions of the model remain inside the 68% CL region for larger values of $\beta$.
@@ I Introduction
The theory of cosmic inflation is now widely accepted as a fundamental component of the standard cosmological model. It was originally proposed to address several critical shortcomings inherent in the Hot Big Bang theory, such as the horizon, flatness, and primordial monopole problems Guth:1981 ; Linde:1982 ; Starobinsky:1980 . Beyond resolving classical puzzles, the inflationary era provides a compelling mechanism for generating primordial density perturbations. These perturbations are originated from quantum fluctuations of the scalar inflaton field Mukhanov:1981 ; Hawking:1982 ; Bardeen:1983 ; Guth:1982pi . After inflation ends, these fluctuations seed the formation of large scale structures and imprint the observed anisotropies in the Cosmic Microwave Background (CMB) radiation.
@@
Among various inflationary frameworks, the $\alpha$-attractor models have received significant attention Kallosh:2013a ; Kallosh:2013b ; Ferrara:2013 ; Ellis:2013 ; Kallosh:2014 ; Rezazadeh2022 . These models are well motivated by supergravity and string theory. Moreover, their predictions for key observables such as the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ were in excellent agreement with earlier observational data Planck:2018inflation ; bk18 . These features have made $\alpha$-attractor models particularly attractive in the context of early universe cosmology. However, recent observational data from ACT DR6, when combined with Planck, DESI BAO, and BICEP/Keck, suggest a slightly higher value for the scalar spectral index $n_{s}$ compared to that inferred from earlier data, namely $n_{s}=0.974\pm 0.003$ ACT:DR6params ; ACT:DR6models . Given these updated constraints, it is necessary to reassess well motivated inflationary scenarios Kallosh:2025act ; Aoki:2025 ; Berera:2025 ; Dioguardi:2025 ; Raidal:2025 ; Salvio:2025 ; Antoniadis:2025 ; Kim:2025 ; Drees:2025 ; Haque:2025 . Within this context, the predictions of $\alpha$-attractor models for $N=60$ now lie close to the edge of the $2\sigma$ region, and are excluded for lower allowed values of the $e$-folds number ACT:DR6params ; ACT:DR6models . This discrepancy motivates refinements to the theoretical predictions of $\alpha$-attractor models and encourages the exploration of additional physical effects that may reconcile these models with the latest cosmological observations.
@@
The presence of a minimal length implies that the standard Heisenberg Uncertainty Principle (HUP) extended to a Generalized Uncertainty Principle (GUP) Luo:2025 ; Bitaj:2024 ; EPL:2023 ; Du:2022 ; Feng:2022 ; Luciano:2021 ; Mohammadi:2017 ; Medved:2004 ; Ali:2012 ; Awad:2014 ; Zhu:2009 ; Das:2008 ; Kempf:1995 . A widely studied GUP model incorporates a quadratic correction in momentum, which is given by
@@ II $\alpha$-attractor inflation modified by GUP
It has been proven that the Friedmann equations are modified in the presence of a minimal length Bitaj:2024 ; Mohammadi:2017 ; Awad:2014 . The modified Friedmann equations in the presence of GUP can be derived from thermodynamical considerations as follows Bitaj:2024
@@
Moreover, the equation of motion of the scalar field is derived by varying the action (3) with respect to $\phi$ as follows
@@
where $({}^{\prime})$ denotes the derivative with respect to $\phi$. It should be noted that within the framework of GUP, the form of the equation of motion (11) is same as that one in the standard model of inflation.
@@
The value of the scalar power spectrum at the pivot scale $k_{\ast}=0.05~{}{\rm Mpc^{-1}}$ has been determined by Planck measurements as ${\cal P}_{s}(k_{\ast})\simeq 2.1\times 10^{-9}$ Planck:2018inflation ; bk18 . Considering Eqs. (13) and (14), the scalar and tensor spectral indexes in Eq. (2) can be estimated as
@@
The observational constraint on $r$ from the latest data of Planck and BICEP/ Keck 2018 is $r<0.036$. In the single field inflationary models, the slow-roll conditions are satisfied when $\dot{\phi}^{2}/2\ll V(\phi)$ and $|\ddot{\phi}|\ll|3H\dot{\phi}|$. Thereunder the equation of motion (11) simplifies to
@@
where $\Lambda$ is the energy scale of the model with dimension of mass, and it can be defined through fixing the amplitude of scalar perturbations at the pivot scale. Furthermore, $\alpha$ is a dimensionless parameter characteristic of the model.
@@ III.1 Behavior of background quantities with GUP corrections
We begin by analyzing the variation of the scalar field at the pivot scale $\phi_{*}$ as a function of GUP parameter $\beta$, as illustrated in Fig. 1. This analysis considers four different values of $\alpha$. Figure 1 shows that i) for a given $\beta$, the value of $\phi_{*}$ increases when the parameter $\alpha$ increases. ii) For a given $\alpha$, the value of $\phi_{*}$ shows a decreasing trend with increasing values of the GUP parameter $\beta$. This indicates that the presence of a minimal length changes the initial value of the inflaton, which can have a significant impact on the model predictions. (iii) For higher values of $\beta$ greater than of order of $O(10^{16})$, the scalar field at the pivot scale $\phi_{*}$ becomes independent of the parameter $\alpha$. Note that throughout the paper, we set $N_{*}=60$ for the horizon exit of the pivot scale. Additionally, all parameters involving mass are normalized by the reduced Planck mass $M_{\rm p}$.
Figure 1: Variation of the scalar field at the pivot scale $\phi_{*}/M_{\rm p}$ versus the GUP parameter $\beta$ for different values of $\alpha$. Here, $N_{*}=60$ is assumed for the horizon exit of the pivot scale.
@@
Figure 2: Evolution of the scalar field $\phi/M_{p}$ as a function of $e$-folds number $N$ for (a) $\alpha=1$ and (b) $\alpha=10$. Each panel displays six curves corresponding to different values of the GUP parameter $\beta=(0,10^{10},10^{11},10^{12},10^{13},10^{14})$. The case of $\beta=0$ refers to the standard model of inflation.
@@
Figure 3: Evolution of the Hubble parameter $H$ as a function of $e$-folds number $N$ for (a) $\alpha=1$ and (b) $\alpha=10$. Each panel displays six curves corresponding to different values of the GUP parameter $\beta=(0,10^{10},10^{11},10^{12},10^{13},10^{14})$.
@@ III.2 Inflationary observables $n_{s}$ and $r$ in the GUP framework
To further investigate the viable parameter space, Fig. 5 presents a phase-space diagram delineating the allowed regions in the $\alpha\,–\,\beta$ plane, based on the 68% and 95% CL from the combined P-ACT-LB-BK18 dataset. The corresponding numerical values extracted from the figure are summarized in Table 1. The analysis indicates that the maximum value of $\alpha$ for which the model predictions lie within the 68% CL contour is approximately $\alpha\simeq 0.74$, which corresponds to a GUP parameter value of $\beta\simeq 1.4\times 10^{13}$, as reported in Table 1. Moreover, the results demonstrate that for higher values of $\beta$, the model predictions remain consistent with the 68% CL region only for comparatively lower values of $\alpha$. Table 1 further shows that as the magnitude of $\beta$ increases, the allowed range of $\alpha$ becomes increasingly narrow. These findings strongly suggest that incorporating GUP effects can substantially improve the observational compatibility of $\alpha$-attractor inflationary model.
@@
$\beta$ $\alpha$ (95% CL) $\alpha$ (68% CL)
[$0,10^{11}$] $\alpha\leq 10$
$10^{13}$ $\alpha\leq 1.50$
@@ IV Conclusions
We concluded that incorporating GUP corrections into the $\alpha$-attractor T-model leads to a markedly improved agreement with recent observational data. In particular, while the standard scenario with $\beta=0$ places the model predictions for $N_{*}=60$ near the edge of the $2\sigma$ region, the presence of minimal length effects for $\beta\gtrsim 10^{13}$ shifts the predicted values of $n_{s}$ and $r$ into the 68% confidence level region favored by the P-ACT-LB-BK18 dataset. Notably, a maximum value of $\alpha\simeq 0.74$ remains compatible with the 68% confidence region at $\beta\simeq 1.4\times 10^{13}$. Furthermore, smaller values of $\alpha$ are found to be viable at larger $\beta$, although the allowed interval for $\alpha$ becomes increasingly narrow as $\beta$ rises. These findings underscore the potential of GUP-induced modifications to enhance the observational viability of $\alpha$-attractor model of inflation.

### v2

@@ Abstract
Here, the $\alpha$-attractor inflation is investigated within a framework incorporating a minimal measurable length, as implemented by the Generalized Uncertainty Principle (GUP). The GUP modifications to the Friedmann equations and cosmological perturbation parameters are employed to assess the model observational viability against the Atacama Cosmology Telescope (ACT) data. Our results indicate that in the $r-n_{s}$ plane, the predictions of the standard $\alpha$-attractor model ($\beta=0$) lies near the $2\sigma$ boundary of joint observations. More interestingly enough is that in the presence of GUP effect, the predictions of the model for the GUP parameter $\beta\gtrsim O(10^{13})$ shifts into the $68\%$ CL interval. This value for $\beta$ is in well agreement with upper bounds on the GUP parameter deduced from cosmological analysis as well as quantum and gravitational experiments.
@@ I Introduction
The theory of cosmic inflation is now widely accepted as a fundamental component of the standard cosmological model. It was originally proposed to address several critical shortcomings inherent in the Hot Big Bang theory, such as the horizon, flatness, and primordial monopole problems Guth:1981 ; Linde:1982 ; Starobinsky:1980 . Beyond resolving classical puzzles, the inflationary era provides a compelling mechanism for generating primordial density perturbations. These perturbations are originated from quantum fluctuations of the scalar inflaton field Mukhanov:1981 ; Hawking:1982 ; Bardeen:1983 ; Guth:1982pi ; Starobinsky:1982ee . After inflation ends, these fluctuations seed the formation of large scale structures and imprint the observed anisotropies in the Cosmic Microwave Background (CMB) radiation.
@@
Among various inflationary frameworks, the $\alpha$-attractor models have received significant attention Kallosh:2013a ; Kallosh:2013b ; Ferrara:2013 ; Ellis:2013 ; Kallosh:2014 ; Rezazadeh:2022 ; Teimoori:2021 ; Carrasco:2015uma . These models are well motivated by supergravity and string theory Kallosh:2013b ; Ferrara:2013 ; Kallosh:2014 ; Carrasco:2015uma . Moreover, their predictions for key observables such as the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ were in excellent agreement with earlier observational data Planck:2018inflation ; bk:18 . These features have made $\alpha$-attractor models particularly attractive in the context of early universe cosmology. However, recent observational data from ACT Data Release 6 (ACT DR6), when combined with Planck, DESI BAO, and BICEP/Keck, suggest a slightly higher value for the scalar spectral index $n_{s}$ compared to that inferred from earlier data, namely $n_{s}=0.974\pm 0.003$ ACT:DR6params ; ACT:DR6models . Given these updated constraints, it is necessary to reassess well motivated inflationary scenarios Kallosh:2025act ; Aoki:2025 ; Berera:2025 ; Dioguardi:2025 ; Raidal:2025 ; Salvio:2025 ; Antoniadis:2025 ; Kim:2025 ; Drees:2025 ; Haque:2025 ; Peng:2025bws ; Liu:2025qca ; Gialamas:2025ofz ; He:2025bli ; Wolf:2025ecy ; Dioguardi:2025mpp . Within this context, the predictions of $\alpha$-attractor models for $N=60$ now lie close to the edge of the $2\sigma$ region, and are excluded for lower allowed values of the $e$-folds number ACT:DR6params ; ACT:DR6models . This discrepancy motivates refinements to the theoretical predictions of $\alpha$-attractor models and encourages the exploration of additional physical effects that may reconcile these models with the latest cosmological observations.
@@
The presence of a minimal length implies that the standard Heisenberg Uncertainty Principle (HUP) extended to a Generalized Uncertainty Principle (GUP), a concept first introduced in the pioneering work Kempf:1995 . For a comprehensive overview, see the review Bosso:2023aht and references therein. Various GUP models and their phenomenological implications have been extensively explored in the literature Luo:2025 ; Bitaj:2024 ; EPL:2023 ; Du:2022 ; Feng:2022 ; Luciano:2021 ; Mohammadi:2017 ; Medved:2004 ; Ali:2012 ; Awad:2014 ; Zhu:2009 ; Das:2008 ; Gialamas:2024ivq ; Segreto:2024vtu ; Banerjee:2022slh ; Feng:2020ams ; Bambi:2007ty ; Barca:2023epu . A widely studied GUP model incorporates a quadratic correction in momentum, which is given by
@@ II $\alpha$-attractor inflation modified by GUP
It has been proven that the Friedmann equations are modified in the presence of a minimal length Bitaj:2024 ; Mohammadi:2017 ; Awad:2014 . The modified Friedmann equations can be derived from thermodynamical considerations. This approach is based on applying the first law of thermodynamics, $dE=TdS$, to the apparent horizon of the universe. The GUP introduces a correction to the standard Bekenstein-Hawking entropy and consequently the standard entropy-area relation is modified. This GUP-corrected entropy, when used in the first law, results in modified Friedmann equations. For a detailed, step-by-step derivation of this procedure, we refer the reader to the work by Awad and Ali Awad:2014 . The modified Friedmann equations are given as
@@
Moreover, the equation of motion for the scalar field $\phi$ is derived from the energy conservation law, $\dot{\rho}_{\phi}+3H(\rho_{\phi}+p_{\phi})=0$. By substituting Eqs. (7) and (8) in the conservation law equation the equation of motion of the scalar field is derived as follows
@@
where $({}^{\prime})$ denotes the derivative with respect to $\phi$. It is crucial to note that while the form of this equation is standard, the GUP effect is included implicitly. The Hubble parameter, $H$, which appears in Eq. (11) is determined by the GUP-modified Friedmann equations. Therefore, the evolution of the scalar field $\phi$ is inherently sensitive to the minimal length corrections through the modified background dynamics.
@@
The value of the scalar power spectrum at the pivot scale $k_{\ast}=0.05~{\rm Mpc^{-1}}$ has been determined by Planck measurements as ${\cal P}_{s}(k_{\ast})\simeq 2.1\times 10^{-9}$ Planck:2018inflation ; bk:18 . Considering Eqs. (13) and (14), the scalar and tensor spectral indexes in Eq. (2) can be estimated as
@@
The observational constraint on $r$ from the latest data of Planck and BICEP/ Keck 2018 is $r<0.036$. In the single field inflationary models, the slow-roll conditions are satisfied when $\dot{\phi}^{2}/2\ll V(\phi)$ and $|\ddot{\phi}|\ll|3H\dot{\phi}|$. There under, the equation of motion (11) simplifies to
@@
where $\Lambda$ is the energy scale of the model with dimension of mass, and it can be defined through fixing the amplitude of scalar perturbations at the pivot scale. Furthermore, $\alpha$ is a dimensionless parameter characteristic of the model. It is worth noting that we focus on the T-model potential as a representative example of the $\alpha$-attractor class. As our results will show, consistency with observational data in the presence of GUP corrections requires large values of the parameter $\beta$, which in turn favors the small-$\alpha$ regime. In this limit, the closely related E-models yield nearly identical predictions for $n_{s}$ and $r$. Therefore, an analysis of the E-model would not alter the main conclusions of this work. For clarity, we limit our scope to the T-model.
@@ III.1 Behavior of background quantities with GUP corrections
We begin by analyzing the variation of the scalar field at the pivot scale $\phi_{*}$ as a function of GUP parameter $\beta$, as illustrated in Fig. 1. This analysis considers four different values of $\alpha$. Figure 1 shows that i) for a given $\beta$, the value of $\phi_{*}$ increases when the parameter $\alpha$ increases. This indicates that the presence of a minimal length changes the initial value of the inflaton, which can have a significant impact on the model predictions. ii) For a given $\alpha$, the value of $\phi_{*}$ decreases as the GUP parameter $\beta$ increases. This reduction is a key advantage of the model with GUP corrections. For large values of $\beta$, the inflaton field $\phi_{*}$ is driven into the sub-Planckian domain. This ensures the inflationary dynamics operate within a valid perturbative regime, which greatly enhances the theoretical self-consistency of the model. iii) For $\beta\gtrsim 10^{18}$, the scalar field at the pivot scale $\phi_{*}$ becomes independent of the parameter $\alpha$. Note that throughout the paper, we set $N_{*}=60$ for the horizon exit of the pivot scale. Additionally, all parameters involving mass are normalized by the reduced Planck mass $M_{\rm p}$.
Figure 1: Field value of the scalar field at the pivot scale $\phi_{*}/M_{\rm p}$ versus the GUP parameter $\beta$ for different values of $\alpha$. Here, $N_{*}=60$ is assumed for the horizon exit of the pivot scale.
@@
Figure 2: Evolution of the scalar field $\phi/M_{p}$ as a function of $e$-folds number NN for (a) $\alpha=1$ and (b) $\alpha=10$. Each panel displays six curves corresponding to different values of the GUP parameter $\beta=(0,10^{10},10^{11},10^{12},10^{13},10^{14})$. The case of $\beta=0$ refers to the standard model of inflation.
@@
Figure 3: Evolution of the Hubble parameter $H$ as a function of $e$-folds number NN for (a) $\alpha=1$ and (b) $\alpha=10$. Each panel displays six curves corresponding to different values of the GUP parameter $\beta=(0,10^{10},10^{11},10^{12},10^{13},10^{14})$.
@@ III.2 Inflationary observables $n_{s}$ and $r$ in the GUP framework
To better understand the direct impact of the GUP parameter on the inflationary observables, we plot the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ as a function of $\beta$ for different values of $\alpha$ (Fig. 5). Figure 5(a) shows the behavior of $n_{s}$ versus $\beta$. The plot reveals a non-monotonic relationship: for each value of $\alpha$, as $\beta$ increases, $n_{s}$ first rises from its standard value, reaches a peak within the observationally favored region, and then decreases. This highlights that for any given $\alpha$, GUP corrections can successfully shift $n_{s}$ into the 68% or 95% CL regions, but only for a specific range of $\beta$ values. Figure 5(b) displays the tensor-to-scalar ratio $r$ as a function of $\beta$. In contrast to the scalar spectral index, $r$ shows a monotonic increase, rising sharply from a small value to a higher plateau as $\beta$ becomes large. This plot also indicates that as the required value of $\beta$ increases, smaller values of $\alpha$ are necessary for the model to remain compatible with the observational bound on $r$.
Figure 5: The inflationary observables as a function of the GUP parameter $\beta$ for several fixed values of $\alpha$. (a) The scalar spectral index $n_{s}$ versus $\beta$. (b) The tensor-to-scalar ratio $r$ versus $\beta$ on a logarithmic scale. The dark and light purple shaded regions show the 68% and 95% CL constraints from the combined P-ACT-LB-BK18 dataset, respectively.
To further investigate the viable parameter space, Fig. 6 presents a phase-space diagram delineating the allowed regions in the $\alpha$$\beta$ plane, based on the 68% and 95% CL from the combined P-ACT-LB-BK18 dataset. The corresponding numerical values extracted from the figure are summarized in Table 1. The analysis indicates that the maximum value of $\alpha$ for which the model predictions lie within the 68% CL contour is approximately $\alpha\simeq 0.74$, which corresponds to the GUP parameter $\beta\simeq 1.4\times 10^{13}$. Note that this value for $\beta$ is in well agreement with upper bounds on the GUP parameter deduced from cosmological analysis as well as quantum and gravitational experiments Luciano:2021 . Moreover, the results demonstrate that for higher values of $\beta$, the model predictions remain consistent with the 68% CL region only for comparatively lower values of $\alpha$. Table 1 further shows that as the magnitude of $\beta$ increases, the allowed range of $\alpha$ becomes increasingly narrow. These findings strongly suggest that incorporating GUP effects can substantially improve the observational compatibility of $\alpha$-attractor inflationary model.
@@
$\beta$ $\alpha~(95\%$ CL) $\alpha~(68\%$CL)
$0,10^{11}]$ $\alpha\leq 10$ -
$10^{13}$ $\alpha\leq 1.50$ -
@@ IV Conclusion
We concluded that incorporating GUP corrections into the $\alpha$-attractor T-model leads to a markedly improved agreement with recent observational data. In particular, while the standard scenario with $\beta=0$ places the model predictions for $N_{*}=60$ near the edge of the $2\sigma$ region, the presence of minimal length effects for $\beta\gtrsim 10^{13}$ shifts the predicted values of $n_{s}$ and $r$ into the 68% confidence level region favored by the P-ACT-LB-BK18 dataset. A key outcome of our analysis is that the GUP parameter $\beta$ directly impacts the inflationary observables. We found that increasing $\beta$ leads to a simultaneous increase in both the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. This effect is crucial, as it shifts the model predictions into closer agreement with the latest observational data.
Notably, a maximum value of $\alpha\simeq 0.74$ remains compatible with the 68% confidence region at $\beta\simeq 1.4\times 10^{13}$. Furthermore, smaller values of $\alpha$ are found to be viable at larger $\beta$, although the allowed interval for $\alpha$ becomes increasingly narrow as $\beta$ rises. These findings underscore the potential of GUP-induced modifications to enhance the observational viability of $\alpha$-attractor model of inflation.
Acknowledgements
The authors thank the referee for his/her valuable comments.